# Ordering in statistical systems on the way to the thermodynamic limit

### v1

@@ 1 Introduction
In order to avoid misunderstanding, let us specify what is the main picture we keep in mind and what are the secondary details that can be neglected and that we will not touch. We consider a system of $N$ particles in volume $V$, hence density $\rho=N/V$, under fixed thermodynamic and external parameters. According to the ideology of the thermodynamic limiting procedure, we then consider the system of a larger number of the same kind of particles in a larger volume, such that the density remains constant and all thermodynamic parameters are the same, being fixed. In that way, we consider not a single system, with increasing volume, but a sequence of systems with increasing sizes keeping the same fixed density and thermodynamic parameters. This is exactly the procedure accepted in the thermodynamic limit.
We consider statistical systems with large numbers of particles $N\gg 1$. Because of this inequality, although the number $N$ can be finite, but, due to its large value, boundary effects can be neglected, especially when $N\rightarrow\infty$. Therefore we do not need to resort to the thermodynamics of small systems of the size of single molecules [8]. Since, we increase the system size explicitly, we do not need to involve finite-size scaling [9] that is used in numerical calculations, when one tries to extrapolate to an infinite system size the measurements accomplished for small systems.
@@
The guiding idea for developing a mathematical approach that could describe the growing order under the increasing system size can be qualitatively characterized as follows. Order, emerging in a system, depends on the relation between an effective correlation length and the system size. The correlation length is defined by the corresponding correlation function. For example, a particular important case of correlation functions is represented by reduced density matrices [22]. The latter can be treated as matrix elements of the reduced density operators $\hat{\rho}_{n}$, where $n=1,2,\ldots$. The norm of a density operator, $\hat{\rho}_{n}||$, is connected with its largest eigenvalue, which, in turn, is related to the effective correlation length. The trace of the density operator is a function of the number of particles $N$. For instance, $\rm{Tr}\rho_{1}=N$. Then the exponent $\omega$, characterizing the relation between the norm and trace of a density operator, describes the dependence of the density operator spectrum on the number of particles, $\hat{\rho}_{1}||=N^{\omega}$. The parameter $\omega$ is an order index showing how quickly the order grows with the increase of the system size. When the order index is small, $\omega\ll 1$, this implies that the correlation length is much shorter that the system effective length, hence there is no long-range order. But when the order index is close to one, then the correlation length becomes comparable to the system effective size, which means that long-range order is getting established in the system.
It is the aim of the present paper to formulate these ideas, employing reduced density matrices for finite systems, in a rigorous quantitative way, to extend the consideration to general correlation operators, and to illustrate the approach by several concrete cases of arising long-range order in the process of growing system size tending from finite $N$ to the thermodynamic limit. We consider the cases of Bose-Einstein condensation, superconducting transition, magnetic transition, and crystal-liquid transition, which, to make the ideas transparent, are treated in mean-field approximation.
@@ 3 Reduced density operators
As is seen, the order index shows how quickly the largest occupation number increases under the growing system size comprising $N$ particles,
@@
It is important to emphasize that the order indices for reduced density operators of any order are well defined for an arbitrary statistical system with a finite number of particles $N$. This follows from the known properties of the reduced density matrices [22], for which the traces and norms of the related reduced density operators are finite for a finite number $N$. This also directly follows from the fact that the density operators are semi-positive trace-class operators for which the inequality (6) is valid.
@@ 5 Bose-Einstein condensation
The order does not appear suddenly, but it grows with increasing system size. The order index (64) is calculated numerically using the solutions to the system of equations (57), (58), and (59). The behavior of the order index for different gas parameters is presented in Fig. 2 as a function of $\ln N$. The increase of order depends on the value of the gas parameter. Thus for $\gamma=0.1$ there is no order for small $N$, where the order index is negative, and the order quickly grows with $N$. For $\gamma=0.5$, there is no order before $N=9$, and the order index reaches $1/2$ close to $N=77$.
@@
while the behavior of the second-order index for finite values of $N$ is shown in Fig. 3. Again, we see that the growth of the order index depends on the interaction strength. Strong interactions make the development of order, under the increasing system size, more difficult.
@@ 7 Magnetic transition
as soon as $M>0$. The behavior of the order index (116) for arbitrary $N$ is shown in Fig. 5.
@@ 8 Crystal-liquid transition
provided that $D>0$. The overall behavior of the order index (138) with respect to the varying particle number $N$ is analogous to that of the magnetic order index (116).
@@ 9 Conclusion
The process describing how the order grows with the increasing system size on the way to the thermodynamic limit is analyzed. For this purpose, a sequence of systems, with fixed parameters, but increasing system size, such that the number of particles $N$ and volume $V$ increase, while the density $\rho=N/V$ kept fixed, is considered. Each system in the sequence is assumed to be an equilibrium statistical system, which requires that the number of its constituents be sufficiently large, $N\gg 1$, and the system lifetime be sufficiently long, longer than the observation time. In this setup, no nonequilibrium effects need to be considered.
The arising and growing order, increasing together with the system size, can be quantified by the order indices that can be introduced for arbitrary trace-class operators, such as the reduced density operators, or other correlation operators. Of course, for quite small systems there could exist finite-size corrections. However, in small systems, where the order has not yet been developed, the order index is also quite small. Noticeable order emerges in sufficiently large systems with the number of particles $N\gg 1$, when boundary corrections do not play important role, and practically disappear for very large $N$, completely vanishing in the thermodynamic limit when $N\rightarrow\infty$.
With the increasing system size, hence the increasing number of particles $N$, the order index grows from zero or even negative, showing the absence of order, to finite positive values, reaching, in the thermodynamic limit, $N\rightarrow\infty$ the maximal possible value. Thus, for the cases where the order is due to some arrangement of individual variables, the order index reaches the value one, while in the case, where the order is caused by pairing correlations, the single-particle order index shows no order and the second-order index can reach the value $1/2$. The order indices, increasing with the growing system size, are considered for the processes of Bose-Einstein condensation, superconductivity, magnetization, and crystallization.
@@
In a microscopic system of just a few particles there can be no order. When the system increases reaching mesoscopic size, where the number of particles $N$ is sufficiently large, there can appear a metastable ordered phase having a finite lifetime. More detailed discussion of characteristic times is postponed to the Appendix. Here we concentrate on spatial correlations that are in the origin of arising order.
@@
The typical lifetime $t_{esc}$, is estimated by the formula (169), where $t_{0}$ is of the order of local equilibration time of $10^{-13}$ s and $c$ is a constant of order one, characterizing the system. As is seen, for the number of particles in the ordered phase $N_{0}$ of order one, the lifetime does not differ much from the local equilibration time, while for $N_{0}$ of order $10$, the lifetime can be of order $10^{-9}$ s. For the number of particles in the system of order $50$, the lifetime of the ordered phase reaches the order of $16$ years. With increasing $N$, the lifetime quickly grows to huge values. This explains why phase transitions and ordered phases that, strictly speaking, in finite systems are metastable, do exist and can be observed without any problem.

### v2

@@ 1 Introduction
In order to avoid misunderstanding, let us specify what is the main picture we keep in mind and what are the secondary details that can be neglected and that we will not touch. We consider a system of NN particles in volume $V$, hence density $\rho=N/V$, under fixed thermodynamic and external parameters. According to the ideology of the thermodynamic limiting procedure, we then consider the system of a larger number of the same kind of particles in a larger volume, such that the density remains constant and all thermodynamic parameters are the same, being fixed. In that way, we consider not a single system, with increasing volume, but a sequence of systems with increasing sizes keeping the same fixed density and thermodynamic parameters. This is exactly the procedure accepted in the thermodynamic limit.
We consider statistical systems with large numbers of particles $N\gg 1$. Because of this inequality, although the number NN can be finite, but, due to its large value, boundary effects can be neglected, especially when $N\rightarrow\infty$. Therefore we do not need to resort to the thermodynamics of small systems of the size of single molecules [8]. Since, we increase the system size explicitly, we do not need to involve finite-size scaling [9] that is used in numerical calculations, when one tries to extrapolate to an infinite system size the measurements accomplished for small systems.
@@
The guiding idea for developing a mathematical approach that could describe the growing order under the increasing system size can be qualitatively characterized as follows. Order, emerging in a system, depends on the relation between an effective correlation length and the system size. The correlation length is defined by the corresponding correlation function. For example, a particular important case of correlation functions is represented by reduced density matrices [22]. The latter can be treated as matrix elements of the reduced density operators $\hat{\rho}_{n}$, where $n=1,2,\ldots$. The norm of a density operator, $\hat{\rho}_{n}||$, is connected with its largest eigenvalue, which, in turn, is related to the effective correlation length. The trace of the density operator is a function of the number of particles NN. For instance, $\rm{Tr}\rho_{1}=N$. Then the exponent $\omega$, characterizing the relation between the norm and trace of a density operator, describes the dependence of the density operator spectrum on the number of particles, $\hat{\rho}_{1}||=N^{\omega}$. The parameter $\omega$ is an order index showing how quickly the order grows with the increase of the system size. When the order index is small, $\omega\ll 1$, this implies that the correlation length is much shorter that the system effective length, hence there is no long-range order. But when the order index is close to one, then the correlation length becomes comparable to the system effective size, which means that long-range order is getting established in the system.
It is the aim of the present paper to formulate these ideas, employing reduced density matrices for finite systems, in a rigorous quantitative way, to extend the consideration to general correlation operators, and to illustrate the approach by several concrete cases of arising long-range order in the process of growing system size tending from finite NN to the thermodynamic limit. We consider the cases of Bose-Einstein condensation, superconducting transition, magnetic transition, and crystal-liquid transition, which, to make the ideas transparent, are treated in mean-field approximation.
@@ 3 Reduced density operators
As is seen, the order index shows how quickly the largest occupation number increases under the growing system size comprising NN particles,
@@
It is important to emphasize that the order indices for reduced density operators of any order are well defined for an arbitrary statistical system with a finite number of particles NN. This follows from the known properties of the reduced density matrices [22], for which the traces and norms of the related reduced density operators are finite for a finite number NN. This also directly follows from the fact that the density operators are semi-positive trace-class operators for which the inequality (6) is valid.
@@ 5 Bose-Einstein condensation
The order does not appear suddenly, but it grows with increasing system size. The order index (64) is calculated numerically using the solutions to the system of equations (57), (58), and (59). The behavior of the order index for different gas parameters is presented in Fig. 2 as a function of $\ln N$. The increase of order depends on the value of the gas parameter. Thus for $\gamma=0.1$ there is no order for small NN, where the order index is negative, and the order quickly grows with NN. For $\gamma=0.5$, there is no order before $N=9$, and the order index reaches $1/2$ close to $N=77$.
@@
while the behavior of the second-order index for finite values of NN is shown in Fig. 3. Again, we see that the growth of the order index depends on the interaction strength. Strong interactions make the development of order, under the increasing system size, more difficult.
@@ 7 Magnetic transition
as soon as $M>0$. The behavior of the order index (116) for arbitrary NN is shown in Fig. 5.
@@ 8 Crystal-liquid transition
provided that $D>0$. The overall behavior of the order index (138) with respect to the varying particle number NN is analogous to that of the magnetic order index (116).
@@ 9 Conclusion
The process describing how the order grows with the increasing system size on the way to the thermodynamic limit is analyzed. For this purpose, a sequence of systems, with fixed parameters, but increasing system size, such that the number of particles NN and volume $V$ increase, while the density $\rho=N/V$ kept fixed, is considered. Each system in the sequence is assumed to be an equilibrium statistical system, which requires that the number of its constituents be sufficiently large, $N\gg 1$, and the system lifetime be sufficiently long, longer than the observation time. In this setup, no nonequilibrium effects need to be considered.
The arising and growing order, increasing together with the system size, can be quantified by the order indices that can be introduced for arbitrary trace-class operators, such as the reduced density operators, or other correlation operators. Of course, for quite small systems there could exist finite-size corrections. However, in small systems, where the order has not yet been developed, the order index is also quite small. Noticeable order emerges in sufficiently large systems with the number of particles $N\gg 1$, when boundary corrections do not play important role, and practically disappear for very large NN, completely vanishing in the thermodynamic limit when $N\rightarrow\infty$.
With the increasing system size, hence the increasing number of particles NN, the order index grows from zero or even negative, showing the absence of order, to finite positive values, reaching, in the thermodynamic limit, $N\rightarrow\infty$ the maximal possible value. Thus, for the cases where the order is due to some arrangement of individual variables, the order index reaches the value one, while in the case, where the order is caused by pairing correlations, the single-particle order index shows no order and the second-order index can reach the value $1/2$. The order indices, increasing with the growing system size, are considered for the processes of Bose-Einstein condensation, superconductivity, magnetization, and crystallization.
@@
In a microscopic system of just a few particles there can be no order. When the system increases reaching mesoscopic size, where the number of particles NN is sufficiently large, there can appear a metastable ordered phase having a finite lifetime. More detailed discussion of characteristic times is postponed to the Appendix. Here we concentrate on spatial correlations that are in the origin of arising order.
@@
The typical lifetime $t_{esc}$, is estimated by the formula (169), where $t_{0}$ is of the order of local equilibration time of $10^{-13}$ s and $c$ is a constant of order one, characterizing the system. As is seen, for the number of particles in the ordered phase $N_{0}$ of order one, the lifetime does not differ much from the local equilibration time, while for $N_{0}$ of order $10$, the lifetime can be of order $10^{-9}$ s. For the number of particles in the system of order $50$, the lifetime of the ordered phase reaches the order of $16$ years. With increasing NN, the lifetime quickly grows to huge values. This explains why phase transitions and ordered phases that, strictly speaking, in finite systems are metastable, do exist and can be observed without any problem.